OpenStudy (swaf11):

solve the equation 2 1/3x=7 7/8 a. 3 3/8 b. 3 1/2 c. 3 5/8 d.4 1/4

OpenStudy (swaf11):

plz help!!!!

OpenStudy (anonymous):

Let's solve your equation step-by-step.

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation.

OpenStudy (swaf11):

kk

OpenStudy (anonymous):

7/3x=63/8

OpenStudy (anonymous):

Step 2: Multiply both sides by 3/7.

OpenStudy (anonymous):

(3/7)*(7/3x)=(3/7)*(63/8)

OpenStudy (anonymous):

x=27/8

OpenStudy (anonymous):

Answer: 3 3/8

OpenStudy (anonymous):

So A

OpenStudy (swaf11):

how did u get x